# Mirror Descent Policy Optimisation for Robust Constrained Markov Decision Processes

### v1

@@ 4.4 Sample-based analysis
(a) The proof uses similar derivations as in Theorem 3 of Liu et al., (2021) and then accounts for the performance difference due to transition kernels.
Note that
$\tilde{\mathbf{V}}_{\pi_{k+1},p_{k+1}}^{\alpha}(\rho)$
$=V_{\pi_{k+1},p_{k+1}}(\rho)+\left\langle\lambda_{k}+\eta_{\lambda}\hat{V}_{\pi_{k},p_{k}}^{1:m}(\rho),V_{\pi_{k+1},p_{k+1}}^{1:m}(\rho)\right\rangle+\frac{\alpha}{1-\gamma}B_{d_{\rho}^{\pi_{k+1},p_{k+1}}}(\pi_{k+1},\pi_{k})$
$\leq V_{\pi^{*},p_{k+1}}(\rho)+\left(\frac{\alpha}{1-\gamma}\left(B_{d_{\rho}^{\pi^{*},p_{k+1}}}(\pi^{*},\pi_{k})-B_{d_{\rho}^{\pi^{*},p_{k+1}}}(\pi^{*},\pi_{k+1})\right)+\Theta(\epsilon)\right)\,,$
where the last step follows from setting $\pi=\pi^{*}$ in Lemma 10 and noting that, since $\lambda_{k,j}+\eta_{\lambda}\hat{V}_{\pi_{k},p_{k}}^{j}(\rho)\geq 0$ by property 2 of Lemma 12 and $V_{\pi^{*},p}^{j}(\rho)\leq 0$ for any $j\in[m]$, their inner product will be negative.
It then follows that
$V_{\pi_{k+1},p_{k+1}}(\rho)-V_{\pi^{*},p_{k}}(\rho)\leq\frac{\alpha}{1-\gamma}\left(B_{d_{\rho}^{\pi^{*},p_{k+1}}}(\pi^{*},\pi_{k+1})-B_{d_{\rho}^{\pi^{*},p_{k+1}}}(\pi^{*},\pi_{k+1})\right)+\Theta(\epsilon)$
$\hskip 142.26378pt-\left\langle\lambda_{k}+\eta_{\lambda}\hat{V}_{\pi_{k},p_{k}}^{1:m}(\rho),V_{\pi_{k+1},p_{k+1}}^{1:m}(\rho)\right\rangle-\frac{\alpha}{1-\gamma}B_{d_{\rho}^{\pi_{k+1},p_{k+1}}}(\pi_{k+1},\pi_{k})\,.$
Filling in the lower bound on the inner product from Eq. 41 from Liu et al., (2021) (see also Lemma 13),
$\left\langle\lambda_{k}+\eta_{\lambda}\hat{V}_{\pi_{k},p_{k}}^{1:m}(\rho),V_{\pi_{k+1},p_{k+1}}^{1:m}(\rho)\right\rangle\geq\frac{1}{2\eta_{\lambda}}\left(\left\lVert\lambda_{k+1}\right\rVert^{2}-\left\lVert\lambda_{k}\right\rVert^{2}\right)+\frac{\eta}{2}\left(\left\lVert V_{\pi_{k},p_{k}}^{1:m}(\rho)\right\rVert^{2}-\left\lVert V_{\pi_{k}+1,p_{k+1}}^{1:m}(\rho)\right\rVert^{2}\right)$
$-2\eta_{\lambda}\left\langle V_{\pi_{k+1},p_{k+1}}^{1:m}(\rho),\epsilon_{k+1}\right\rangle-\frac{\gamma^{2}\eta_{\lambda}}{(1-\gamma)^{4}}B_{d_{\rho}^{\pi_{k+1},p_{k+1}}}(\pi_{k+1},\pi_{k})\,,$
we get
$V_{\pi_{k+1},p_{k+1}}(\rho)-V_{\pi^{*},p_{k+1}}(\rho)\leq(1+\Delta_{p})\left(\frac{\alpha}{1-\gamma}\left(B_{d_{\rho}^{\pi^{*},p_{k}}}(\pi^{*},\pi_{k})-B_{d_{\rho}^{\pi^{*},p_{k}}}(\pi^{*},\pi_{k+1})\right)+\Theta(\epsilon)\right)$
$+\frac{1}{2\eta_{\lambda}}\left(\left\lVert\lambda_{k}\right\rVert^{2}-\left\lVert\lambda_{k+1}\right\rVert^{2}\right)+\frac{\eta}{2}\left(\left\lVert V_{\pi_{k}+1,p_{k+1}}^{1:m}(\rho)\right\rVert^{2}-\left\lVert V_{\pi_{k},p_{k}}^{1:m}(\rho)\right\rVert^{2}\right)$
$+2\eta_{\lambda}\left\langle V_{\pi_{k+1},p_{k+1}}^{i:m}(\rho),\epsilon_{k+1}\right\rangle-\frac{\alpha(1-\gamma)^{3}-\gamma^{2}\eta_{\lambda}}{(1-\gamma)^{4}}B_{d_{\rho}^{\pi_{k+1},p_{k+1}}}(\pi_{k+1},\pi_{k})$
$\leq\frac{\alpha}{1-\gamma}\left(B_{d_{\rho}^{\pi^{*},p_{k}}}(\pi^{*},\pi_{k})-B_{d_{\rho}^{\pi^{*},p_{k}}}(\pi^{*},\pi_{k+1})\right)+\Theta(\epsilon)$
$+\frac{1}{2\eta_{\lambda}}\left(\left\lVert\lambda_{k}\right\rVert^{2}-\left\lVert\lambda_{k+1}\right\rVert^{2}\right)+\frac{\eta}{2}\left(\left\lVert V_{\pi_{k}+1,p_{k+1}}^{1:m}(\rho)\right\rVert^{2}-\left\lVert V_{\pi_{k},p_{k}}^{1:m}(\rho)\right\rVert^{2}\right)$
$+2\eta_{\lambda}\left\langle V_{\pi_{k+1},p_{k+1}}^{i:m}(\rho),\epsilon_{k+1}\right\rangle$
Note that the upper bound on $\eta_{\lambda}/2\left\lVert V_{\pi_{k},p_{k}}^{1:m}(\rho)-V_{\pi_{k+1},p_{k+1}}^{1:m}(\rho)\right\rVert^{2}\leq B_{d_{\rho}^{\pi_{k+1},p_{k+1}}}(\pi_{k+1},\pi_{k})$ is analogous to Eq.16 of Liu et al., (2021) but with a divergence term depending on $p_{k+1}$ rather than $p_{k}$. Note that the term $\frac{\alpha(1-\gamma)^{3}-\gamma^{2}\eta_{\lambda}}{(1-\gamma)^{4}}B_{d_{\rho}^{\pi_{k+1},p_{k+1}}}(\pi_{k+1},\pi_{k})\geq 0$, so it drops from the upper bound in the subsequent analysis.
Due to the approximations and Lemma 13, the regret introduces a term
$\Delta_{k}=\Theta(\epsilon)+\left\langle\lambda_{k-1},\epsilon_{k}\right\rangle-\eta_{\lambda}\left\langle\epsilon_{k-1},V_{\pi_{k},p_{k}}^{1:m}(\rho)\right\rangle+\eta_{\lambda}\left\langle\epsilon_{k}+2V_{\pi_{k},p_{k}}^{1:m}(\rho),\epsilon_{k}\right\rangle\,.$
Using Lemma 21 and telescoping, it follows that
$\leq\sum_{k=1}^{K}\frac{\alpha}{1-\gamma}\left(B_{d_{\rho}^{\pi^{*},p_{k}}}(\pi^{*},\pi_{k})-B_{d_{\rho}^{\pi^{*},p_{k}}}(\pi^{*},\pi_{k+1})\right)+\Delta_{k}$
$+\frac{\eta_{\lambda}}{2}\left(\left\lVert V_{\pi_{k},p_{k}}^{1:m}(\rho)\right\rVert^{2}-\left\lVert V_{\pi_{k-1},p_{k-1}}^{1:m}(\rho)\right\rVert^{2}\right)+\frac{1}{2\eta_{\lambda}}\left(\left\lVert\lambda_{k-1}\right\rVert^{2}-\left\lVert\lambda_{k}\right\rVert^{2}\right)$
$\leq\sum_{k=1}^{K}\frac{\alpha}{1-\gamma}\left(B_{d_{\rho}^{\pi^{*},p_{k}}}(\pi^{*},\pi_{k-1})-B_{d_{\rho}^{\pi^{*},p_{k}}}(\pi^{*},\pi_{k})+(p_{k}-p_{k-1})\log(A)\right)+\Delta_{k}$
[…]
$\leq\frac{\alpha}{1-\gamma}\left(B_{d_{\rho}^{\pi^{*},p_{0}}}(\pi^{*},\pi_{0})-B_{d_{\rho}^{\pi^{*},p_{K}}}(\pi^{*},\pi_{K})\right)+\frac{\eta_{\lambda}}{2}\left(\left\lVert V_{\pi_{K},p_{K}}^{1:m}(\rho)\right\rVert^{2}-\left\lVert V_{\pi_{0},p_{0}}^{1:m}(\rho)\right\rVert^{2}\right)$
$+\frac{1}{2\eta_{\lambda}}\left(\left\lVert\lambda_{0}\right\rVert^{2}-\left\lVert\lambda_{K}\right\rVert^{2}\right)+\alpha\left\lVert p_{K}-p_{0}\right\rVert_{1}\mathcal{O}\left(\frac{\log(A)}{1-\gamma}\right)+\sum_{k=1}^{K}\Delta_{k}$
$\leq\frac{\alpha}{1-\gamma}\left(B_{d_{\rho}^{\pi^{*},p_{0}}}(\pi^{*},\pi_{0})-B_{d_{\rho}^{\pi^{*},p_{K}}}(\pi^{*},\pi_{K})\right)+\frac{\eta_{\lambda}m}{(1-\gamma)^{2}}+\alpha\Delta_{p}\mathcal{O}\left(\frac{\log(A)}{1-\gamma}\right)+\sum_{k=1}^{K}\Delta_{k}\,,$ (44)
where the last step follows from $\left\lVert\lambda_{0}\right\rVert\leq\frac{m\eta_{\lambda}^{2}}{(1-\gamma)^{2}}$ and $\left\lVert V_{\pi_{K},p_{K}}(\rho)\right\rVert^{2}\leq\frac{m}{(1-\gamma)^{2}}$. Since $\alpha=\frac{1-\gamma}{\eta}$, $\eta_{\lambda}=\frac{\alpha}{(1-\gamma)^{3}}{2\gamma^{2}}$, $\frac{\alpha}{1-\gamma}\log(A)+\frac{\eta_{\lambda}m}{(1-\gamma)^{2}}=\mathcal{O}(1)$. Moreover, since $\Delta_{k}=\mathcal{O}(\max\left\{1,\left\lVert\lambda_{k}\right\rVert_{1}\right\}\epsilon)$ and Lemma 22 shows that with probability $1-\delta$, $\left\lVert\lambda_{k}\right\rVert\leq\left\lVert\lambda_{k}\right\rVert_{1}=\mathcal{O}(1)$, we have $\Delta_{k}=\mathcal{O}(1)$ with probability $1-\delta$. Finally, $\alpha\Delta_{p}\mathcal{O}\left(\frac{\log(A)}{1-\gamma}\right)$ is also $\mathcal{O}(1)$ since $\Delta_{p}=\Theta\left(\frac{(1-\gamma)}{\alpha\log(A)}\right)$. All terms then reduce to $\mathcal{O}(\epsilon)$ after division by $K=\Theta(\frac{1}{\epsilon})$.
it follows that
$\frac{1}{K}\sum_{k=1}^{K}V_{\pi_{k},p_{k}}^{j}(\rho)$ $=\frac{1}{K}\sum_{k=1}^{K}\left(\hat{V}_{\pi_{k},p_{k}}^{j}(\rho)-\epsilon_{k,j}\right)$
[…]
$\leq\frac{\left\lVert\lambda^{*}\right\rVert+\left\lVert\lambda^{*}-\lambda_{K}\right\rVert}{\eta_{\lambda}K}-\frac{1}{K}\sum_{k=1}^{K}\epsilon_{k,j}\quad\text{(since $\lambda_{K,j}\leq\left\lVert\lambda_{K}\right\rVert\leq\left\lVert\lambda^{*}\right\rVert_{+}\left\lVert\lambda^{*}-\lambda_{K}\right\rVert$)}$
[…]
$=\mathcal{O}(\epsilon)\quad\text{( since $K=\Theta\left(\frac{1}{\epsilon}\right)$)}\,.$
c) For any $k\in[K]$, it follows from Lemma 2 that
$t_{k}^{\prime}=\Theta\left(\frac{F_{\lambda_{k}}M}{\epsilon_{k}^{\prime}(1-\gamma)}\right)$
is sufficient to obtain an $\epsilon_{k}^{\prime}$-optimal transition kernel from LTMA. Using Lemma 7, combining the errors from a) and b) for the cost and constraint-costs, and noting that $\epsilon_{k}^{\prime}=\Theta(\epsilon)$, it follows that
$\frac{1}{K}\sum_{k=1}^{K}\left(\Phi(\pi_{k})-\Phi(\pi^{*})\right)$ $\leq\frac{1}{K}\sum_{k=1}^{K}\left(\mathbf{V}_{\pi_{k},p_{k}}(\rho)+\epsilon_{k}^{\prime}-\mathbf{V}_{\pi^{*},p_{k}}(\rho)\right)$
$=\mathcal{O}(F_{\lambda}\epsilon+\epsilon)$
$=\mathcal{O}(F_{\lambda}\epsilon)\,,$
where $F_{\lambda}=\max_{k\in[K]}F_{\lambda_{k}}$. Due to the results in Lemma 22, $\lambda_{k}=\mathcal{O}(1)$ and $F_{\lambda}=\mathcal{O}(1)$. Moreover $\eta_{\lambda}V_{\pi_{k},p_{k}}^{j}\leq\frac{1}{1-\gamma}=\mathcal{O}(1)$ for all $j\in[m]$ and all $k\in[K]$, which concludes the proof for both the augmented and unaugmented case.
Sample complexity: Following the loop in Algorithm 1, plugging in the settings from Lemma 9), and omitting logarithmic factors, the total number of calls to the generative model is given by
$T$ $=\sum_{k=1}^{K}\left(M_{V,k}N_{V,k}+\sum_{t=0}^{t_{k}-1}M_{Q,k}N_{Q,k}+t_{k}^{\prime}\right)$
$=\mathcal{O}(\epsilon^{-1})\left(\epsilon^{-2}\tilde{\mathcal{O}(1)}+\tilde{\mathcal{O}}(1)\epsilon^{-2}\tilde{\mathcal{O}}(1)\right)+\sum_{k=1}^{K}t_{k}^{\prime}$
$=\tilde{\mathcal{O}}(\epsilon^{-3}+\epsilon^{-1}\frac{F_{\lambda_{k}}M}{\epsilon^{\prime}(1-\gamma)})$
$=\tilde{\mathcal{O}}(\epsilon^{-3})\,,$
where the last step follows from $\epsilon^{\prime-1}=\mathcal{O}(\epsilon^{-1})$, $F_{\lambda}=\mathcal{O}(1)$, and $M$ is a problem-specific constant that depends on the space of policies and the uncertainty set. ∎
@@ 4.5 KL-based Mirror Descent in continuous state-action spaces
To apply the above analysis in continuous state-action spaces, we follow a similar algorithm but replace the tabular approximation with function approximation (see Algorithm 2). A further change is that we still seek to maintain similar theoretical results, so we motivate the extension to continuous state-action spaces using a continuous pseudo-KL divergence of occupancy. Lemma 10 in Liu et al., (2021) defines a pseudo-KL divergence of occupancy within a discrete state-action space, which is shown to be a Bregman divergence of occupancy distributions. To demonstrate the wider applicability of the theory, we extend this definition straightforwardly to continuous state-action spaces below.
@@
Algorithm 2 Robust Sample-based PMD-PD (continuous setting)
Inputs: Discount factor $\gamma\in[0,1)$, sample sizes $M_{Q,k},M_{V,k}$ and episode lengths $N_{Q,k},N_{V,k}$ for all $k\in[K]$, learning rate $\eta$, dual learning rate $\eta_{\lambda}$, error tolerance for LTMA $\epsilon_{0}^{\prime}>0$, and penalty coefficients $\alpha>0$ and $\alpha_{p}>0$.
Initialise: $\pi_{0}$ as uniform random policy, $\lambda_{0,j}=\max\left\{0,-\eta_{\lambda}\hat{V}_{\pi_{0},p_{0}}(\rho)\right\}$, $p_{0}$ as the nominal transition kernel $\bar{p}$.
for $k\in\left\{1,\dots,K\right\}$ do
     for $t\in\left\{0,\dots,t_{k}-1\right\}$ do
         Generate $M_{Q,k}$ samples of length $N_{Q,k}$ based on $\pi_{k}^{t}$ and $p_{k}$.
         Update $\hat{\mathbf{Q}}_{\pi_{k}^{t},k}\leftarrow\mathrm{arg}\min_{Q\in\mathcal{F}_{Q}}\frac{1}{M_{Q,k}}\sum_{j=1}^{M_{Q,k}}\left(Q(s_{0}^{j},a_{0}^{j})-\sum_{l=0}^{N_{V,k}-1}\gamma^{l}\left[\tilde{\mathbf{c}}_{k}(s_{l}^{j},a_{l}^{j})\right]\right)^{2}$.
         $\triangleright$ Policy mirror descent
         $\theta_{k}^{t+1}\leftarrow\mathrm{arg}\min_{\theta\in\Theta}\eta\left\langle\nabla_{\theta}\hat{\mathbf{Q}}_{\pi_{k}^{t},p_{k}},\theta\right\rangle+\alpha B_{d_{\rho}^{\pi_{\theta},p_{k}}}(\theta,\theta_{k})$. $\triangleright$ separate value and divergence estimates
         Define $\pi_{k}^{t+1}:=\pi_{\theta_{k}^{t+1}}$
     end for
     Define $\pi_{k+1}:=\pi_{k}^{t_{k}}$
     $\triangleright$ Update transition kernel with LTMA (error tolerance fixed to $\epsilon_{0}^{\prime}$ or $\epsilon_{k+1}^{\prime}=\gamma\epsilon_{k}^{\prime}$)
     for $t^{\prime}\in\left\{0,\dots,t_{k}^{\prime}-1\right\}$ do
         $\xi_{k}^{t^{\prime}+1}\leftarrow\mathrm{arg}\max_{\xi\in\mathcal{U}}\eta_{p}\left\langle\nabla_{\xi}\hat{\mathbf{V}}_{\pi_{k}^{t},p_{k}^{t^{\prime}}}(\rho),\xi\right\rangle-\alpha_{p}B_{d_{\rho}^{\pi_{k+1},p_{k}^{t^{\prime}}}}(\xi,\xi_{k}^{t^{\prime}})$.  $\triangleright$ Monte Carlo or other techniques
         Define $p_{k}^{t^{\prime}+1}:=p_{\xi_{k}^{t^{\prime}+1}}$.
     end for
     Define $p_{k+1}:=p_{k}^{t_{k}^{\prime}}$.
     $\triangleright$ Dual update of augmented Lagrangian multipliers
     Generate $M_{V,k+1}$ samples of length $N_{V,k+1}$ based on $\pi_{k+1}$ and $p_{k+1}$.
end for
@@ 4.6 MDPO-Robust-Lag: a practical implementation
based on $t_{k}$ steps of SGD steps per batch $k$, which corresponds to the macro-iteration. For macro-iteration $k$, it defines the policy gradient for $\theta=\theta_{k}^{t}$ at any iteration $t\in[t_{k}]$ as
@@
where in our case the advantage $\hat{A}_{\pi_{\theta_{k}},p_{k}}(s,a)$ is implemented based on the Lagrangian. The use of the importance weight $\frac{\pi_{\theta}}{\pi_{\theta_{k}}}$ corrects for the deviation from the theory, where it is assumed that the data at each epoch is generated from $\pi_{\theta}$ rather than $\pi_{\theta_{k}}$. However, there still remains some gap of the practical implementation to our theoretical algorithm in that the state occupancy of MDPO is given by the old policy whereas in theory it is based on the current. The implementation of advantage values is based on generalised advantage estimation (Schulman et al.,, 2018) and the critic estimates the value and the constraint-costs based on an MLP architecture with $1+m$ heads. The update of the multipliers is based on the observed constraint-costs in separate samples.
To form a robust-constrained variant of MDPO, we use Lagrangian relaxation similar to the above (i.e. Algorithm 1), which differs from Adversarial RCPG (Bossens,, 2024) in that multiple constraints are considered and that we preserve the structure of the PMD-PD algorithm (Liu et al.,, 2021) and the use of LTMA similar follows the TMA implementation from Wang and Petrik, (2024), which applies a projected gradient descent based on clipping to the uncertainty set bounds. Due to its similarity to PPO-Lagrangian (or PPO-Lag for short) (Achiam et al.,, 2017) and the use of updates in the robust Lagrangian, we name the algorithm MDPO-Robust-Lag. While the MDPO-Robust-Lag is based on traditionally clipping the multiplier, we also implement a version with the augmented Lagrangian as suggested by the theory and shown in Algorithm 1 and  2, called MDPO-Robust-Augmented-Lag. The practical implementation of the KL estimate follows the standard implementation in StableBaselines3, i.e. $D_{\text{KL}}(\pi(\cdot|s),\pi^{\prime}(\cdot|s))\approx\mathbb{E}\left[e^{\log(\pi(a|s))-\log(\pi^{\prime}(a|s))}-1-(\log(\pi(a|s))-\log(\pi^{\prime}(a|s)))\right]$.
In traditional optimisation settings, $\alpha_{k}^{-1}$ is decreased across iterations to give a larger penalty at the end of the iterations (Beck and Teboulle,, 2003). However, if the LTMA update is strong and the policy cannot take sufficiently large learning steps, it may sometimes lead to a gradual loss of performance and premature convergence. We especially believe this to be a problem in the context of constrained optimisation since the dual variable, and therefore the objective of LTMA, is often subject to large changes due to constraints becoming active or inactive. To counter this issue in practice, we introduce restarts after every time the LTMA update to $\xi$ was larger than a particular threshold, such that
$\alpha_{k}=\frac{1}{1-(k-t^{\prime})/(K-t^{\prime})}\,,$ (49)
where $t^{\prime}$ indicates the restart time. In the experiments, the restart is done every time any dimension $i$ of the update is larger than half the maximal distance from nominal, i.e. larger than $0.5\max_{\xi_{i}\in\mathcal{U}_{\xi_{i}}}|\xi_{i}-\bar{\xi}_{i}|$.
@@
To demonstrate the use of MDPO for RCMDPs, we assess MDPO-Robust-Lag on three RCMDP domains. We compare MDPO-Robust-Lag to robust-constrained algorithms from related algorithm families, namely Monte Carlo based optimisation (RMCPMD (Wang and Petrik,, 2024) in particular) and proximal policy optimisation with function approximiation (PPO-Lag (Achiam et al.,, 2017) in particular). All the involved algorithm families (MDPO, MC, and PPO) are included with four variants, namely with and without Lagrangian (indicated by the -Lag suffix) and with and without robustness training (indicated by the R- prefix or the -Robust- infix). Additionally, the MDPO-Lag implementations also have Augmented-Lag variants which use the augmented Lagrangian as proposed in the theory section. Further implementation details of the algorithms included to the study can be found in Appendix D.
After training with the above-mentioned algorithms, the agents are subjected to a test which presents transition dynamics based on the distortion levels setup in Wang and Petrik, (2024). The test procedure follows the same general procedure for all domains. For distortion level $x\in[0,1]$, the agent is subjected to a perturbation such that if the nominal model is $\xi_{c}$ and the uncertainty set varies the parameter in $[\underline{\xi},\overline{\xi}]$, the transition dynamics kernel parameter of the test with distortion level $x$ is given for each dimension by either $\xi_{i}=\xi_{c}+x^{2}(\overline{\xi}-\xi_{c})$ or $\xi_{i}=\xi_{c}+x^{2}(\underline{\xi}-\xi_{c})$. We slightly improve the test evaluation by considering all the $\pm$ directions rather than just a single one, amounting to a total number of test evaluations $N=n_{\text{test}}\times 2^{n}$, where $n$ is the dimensionality of the uncertainty set parametrisation and $n_{\text{test}}$ is the number of episodes per perturbation.
@@
We report the results based on two distinct sample budgets, one with a limited number of time steps (200–500 times the maximal number of time steps in the episode) and one with a large number of time steps (2000–5000 times the maximal number of time steps in the episode). This setup allows to assess the sample-efficiency and convergence properties more clearly. Following Eq. 49, we experiment with a restart schedule to set the KL penalty of MDPO-Robust-Lag. As shown in Appendix G, the performance of the restart schedule is generally close to that of the linear schedule across the three included domains. However, in combination with the augmented Lagrangian, we observe several results that far exceed the performance of the linearly decreasing schedule for $\alpha_{k}^{-1}=1-k/K$; therefore, all the results shown below for MDPO-Robust-Lag algorithms are based on the restart schedule. We also compare to the geometrically increasing schedule $\alpha_{k}^{-1}=\alpha_{k-1}^{-1}/\gamma$ proposed in Wang and Petrik, (2024), which overall performs similarly to the restart schedule. However, with regard to the crucial unsigned $R_{\text{pen}}$ score, only MDPO-Robust-Augmented-Lag with restart schedule appears consistently at the top or near the top for all the domains and sample budgets.

### v2

@@ 5.4 Analysis of Robust Sample-based PMD-PD
A proof sketch is given below for convenience while the full deriviation is given in Appendix D.6.2.
(a) The proof uses Lemma 10 and Lemma 14 to obtain a telescoping sum similar to the analysis of Theorem 3 in Liu et al., 2021a . To account for transition kernel changes, it then upper bounds the cumulative regret difference using Lemma 23) and the Bregman divergence terms in the RHS of Eq. 36 using Lemma 24. This then results in a telescoping sum.
(b): The proof uses the update rule of the dual variable and its relation to the approximate constraint-cost to obtain a telescoping sum and an additional term $\mathcal{O}(\epsilon)$.
c): The proof uses Lemma 8 to upper bound the robust Lagrangian regret with respect to $(\pi^{*},p^{*},\lambda^{*})$ by using the regret bounds with respect obtained in a) and b). Noting that the regret from oracle-based LTMA (Lemma 2) is at most $\epsilon_{k}^{\prime}=\mathcal{O}(\epsilon)$ for each $k$, and that the average regret induced by the dual variable is bounded by $\mathcal{O}(\epsilon)$ (via Lemma 28).
d): The proof is analogous to part c) but now applies Approximate TMA using the settings of Lemma 7.
Sample complexity: one may then apply Algorithm 2, plug in the settings from Lemma 9 and Lemma 7 to obtain
$T$ $=\sum_{k=1}^{K}\left(M_{V,k}N_{V,k}+\sum_{t=0}^{t_{k}-1}M_{Q,k}N_{Q,k}+\sum_{t=0}^{t_{k}^{\prime}-1}M_{G,t}N_{G,t}\right)=\tilde{\mathcal{O}}(\epsilon^{-3})\,.$
@@ 5.5 KL-based Mirror Descent in continuous state-action spaces
To generalise the algorithm to continuous state-action spaces, we follow a similar algorithm but replace the tabular approximation with function approximation (see Algorithm 3). A further required change is related to the weighted Bregman divergence term formulated in Eq. 13. While the weighted Bregman divergence is not a proper Bregman divergence over policies, it is over the occupancy distribution, i.e. $B_{d_{\rho}^{\pi,p}}(\pi,\pi^{\prime})=B(d_{\rho}^{\pi,p},d_{\rho}^{\pi^{\prime},p})$, as shown in the pseudo-KL divergence in Lemma 10 of Liu et al., 2021b . To maintain similar theoretical results in continuous state-action spaces, we extend the discrete pseudo-KL divergence of occupancy to a continuous pseudo-KL divergence of occupancy.
@@ 5.6 MDPO-Robust-Lag: a practical implementation
based on $t_{k}$ SGD steps per batch $k$, which corresponds to the macro-iteration. For macro-iteration $k$, it defines the policy gradient for $\theta=\theta_{k}^{t}$ at any iteration $t\in[t_{k}]$ as
@@
where in our case the advantage $\hat{A}_{\pi_{\theta_{k}},p_{k}}(s,a)$ is implemented based on the Lagrangian. The use of the importance weight $\frac{\pi_{\theta}}{\pi_{\theta_{k}}}$ is used in MDPO to correct for the deviation from mirror descent theory, where it is assumed that the data at each epoch is generated from $\pi_{\theta}$ rather than $\pi_{\theta_{k}}$. In line with our theory, our implementation performs the critic updates together with policy updates, which deviates from the original MDPO implementation. There still remains some gap of the practical implementation to our theoretical algorithm in that the state occupancy of MDPO is given by the old policy whereas in theory it is based on the current. The implementation of advantage values is based on generalised advantage estimation (Schulman et al.,, 2018) and the critic estimates the value and the constraint-costs based on an MLP architecture with $1+m$ heads. The update of the multipliers is based on the observed constraint-costs in separate samples.
To form a robust-constrained variant of MDPO, we use Lagrangian relaxation similar to the above (i.e. Algorithm 2), which differs from Adversarial RCPG (Bossens,, 2024) in that multiple constraints are considered and that we preserve the structure of the PMD-PD algorithm (Liu et al., 2021b, ) and the use of LTMA similar follows the TMA implementation from Wang et al., (2024), which applies a projected gradient descent based on clipping to the uncertainty set bounds. Due to its similarity to PPO-Lagrangian (or PPO-Lag for short) (Ray et al.,, 2019) and the use of updates in the robust Lagrangian, we name the algorithm MDPO-Robust-Lag. While the MDPO-Robust-Lag is based on traditionally clipping the multiplier, we also implement a version with the augmented Lagrangian as suggested by the theory and shown in Algorithm 2 and  3, called MDPO-Robust-Augmented-Lag. The practical implementation of the KL estimate follows the standard implementation in StableBaselines3, i.e. $D_{\text{KL}}(\pi(\cdot|s),\pi^{\prime}(\cdot|s))\approx\mathbb{E}\left[e^{\log(\pi(a|s))-\log(\pi^{\prime}(a|s))}-1-(\log(\pi(a|s))-\log(\pi^{\prime}(a|s)))\right]$.
@@
To demonstrate the use of MDPO for RCMDPs, we assess MDPO-Robust-Lag on three RCMDP domains. We compare MDPO-Robust-Lag to robust-constrained algorithms from related algorithm families, namely Monte Carlo based optimisation (RMCPMD (Wang et al.,, 2024) in particular) and proximal policy optimisation with function approximiation (PPO-Lag (Achiam et al.,, 2017) in particular). All the involved algorithm families (MDPO, MC, and PPO) are included with four variants, namely with and without Lagrangian (indicated by the -Lag suffix) and with and without robustness training (indicated by the R- prefix or the -Robust- infix). Additionally, the MDPO-Lag implementations also have Augmented-Lag variants which use the augmented Lagrangian as proposed in the theory section. Further implementation details of the algorithms included to the study can be found in Appendix E.
After training with the above-mentioned algorithms, the agents are subjected to a test which presents transition dynamics based on the distortion levels setup in Wang et al., (2024). The test procedure follows the same general procedure for all domains. For distortion level $x\in[0,1]$, the agent is subjected to a perturbation such that if the nominal model is $\xi_{c}$ and the uncertainty set varies the parameter in $\underline{\xi},\overline{\xi}]$, the transition dynamics kernel parameter of the test with distortion level $x$ is given for each dimension by either $\xi_{i}=\xi_{c}+x^{2}(\overline{\xi}-\xi_{c})$ or $\xi_{i}=\xi_{c}+x^{2}(\underline{\xi}-\xi_{c})$. We slightly improve the test evaluation by considering all the $\pm$ directions rather than just a single one, amounting to a total number of test evaluations $N=n_{\text{test}}\times 2^{n}$, where $n$ is the dimensionality of the uncertainty set parametrisation and $n_{\text{test}}$ is the number of episodes per perturbation.
@@
We report the results based on two distinct sample budgets, one with a limited number of time steps (200–500 times the maximal number of time steps in the episode) and one with a large number of time steps (2000–5000 times the maximal number of time steps in the episode). This setup allows to assess the sample-efficiency and convergence properties more clearly.
An additional set of experiments is to establish which kind of schedule for $\alpha_{k}$ works best for MDPO algorithms, and the robust(-constrained) variants in particular (see Apppendix H). From this set of experiments, it is confirmed that a fixed setting works reasonably well across problems. The best overall performance for robust-constrained RL is obtained by MDPO-Robust-Augmented-Lag using a batch-based linear schedule with restarts, and a geometric schedule achieves high performance with MDPO-Robust-Lag. For simplicity, the fixed schedule is used for all MDPO based algorithms throughout the remaining experiments.